OpenStudy (anonymous):

simplify completely: x^2+x-12/x^2-x-20 times 3x^2-24x+45/12x^2-48 x-60

OpenStudy (anonymous):

\[ \frac{x^{2}+x-12}{x^{2}-x-20} \times \frac{ 3x^{2}-24x+45 }{ 12x^{2}-48x-60 }\]